OpenStudy (anonymous):

what would the monomial of x^3y^2y be?

OpenStudy (anonymous):

Well, what is \(y^2 \cdot y\)

OpenStudy (anonymous):

y^3?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

So \(x^3y^2y\) = ?

OpenStudy (anonymous):

Does it help if I write it as \(x^3\cdot y^2 \cdot y\)

OpenStudy (toxicsugar22):

hi can u help me after her please

OpenStudy (anonymous):

would it be x^3y^3?

OpenStudy (anonymous):

that's a start, but now you can simplify that a bit also

OpenStudy (anonymous):

because \(x^3y^3 = (?)^3\)

OpenStudy (anonymous):

xy

OpenStudy (anonymous):

indeed. so \(x^3y^2y = (xy)^3\)

OpenStudy (anonymous):

ok thanks:)